OpenStudy (anonymous):

solve using interval notation x<-1 or x>1

OpenStudy (saifoo.khan):

0

OpenStudy (anonymous):

Thank you! Man you are quick! LOL

OpenStudy (anonymous):

\[(-\infty,-1)\cup (1,\infty)\]

OpenStudy (anonymous):

yeah quick, but unfortunately not right.

OpenStudy (anonymous):

these are two intervals, so you need to write two intervals.

OpenStudy (anonymous):

unfortunately it says you are both wrong.

OpenStudy (saifoo.khan):

wait, i think u hav to show it on the number line??

OpenStudy (anonymous):

i don't know what "it" is but the first inequality says \[x<-1\] and that is \[(-\infty <-1)\]

OpenStudy (anonymous):

I got that part right but when it asked for the solution both were wrong :(

OpenStudy (anonymous):

second says \[x>1\] and that is \[(1,\infty)\]

OpenStudy (anonymous):

that is "interval" notation. you can bet the ranch on it

OpenStudy (anonymous):

I wrote what you put and it said incorrect.

OpenStudy (anonymous):

what is the "it"?

OpenStudy (saifoo.khan):

i googled the interval notation.. it was like this..

OpenStudy (anonymous):

You were right I typed it in wrong.

OpenStudy (anonymous):

well that is nonsense. it says \[x<-1\] or \[x>1\] and that is not what you have in the picture

OpenStudy (anonymous):

whew!

OpenStudy (anonymous):

thought i was losing my feeble mind

OpenStudy (saifoo.khan):

lol

OpenStudy (anonymous):

LOL. Thank you both! :)

OpenStudy (anonymous):

yw